# Enhancing the hyperpolarizability of crystals with quantum geometry

### v1

@@ Abstract
We demonstrate that higher-order electric susceptibilities in crystals can be enhanced and understood through nontrivial topological invariants and quantum geometry, using one-dimensional $\pi$-conjugated chains as representative model systems. First, we show that the crystalline-symmetry-protected topology of these chains imposes a lower bound on their quantum metric and hyperpolarizabilities. Second, we employ numerical simulations to reveal the tunability of non-linear, quantum geometry-driven optical responses in various one-dimensional crystals in which band topology can be externally controlled. Third, we develop a semiclassical picture to deliver an intuitive understanding of these effects. Our findings offer a firm interpretation of otherwise elusive experimental observations of colossal hyperpolarizabilities and establish guidelines for designing topological materials of any dimensionality with enhanced non-linear optical properties.
@@
Introduction.— Non-linear optics offers a rich platform to explore a broad range of intriguing phenomena, with implications ranging from fundamental science to engineering Boyd (2019). One-dimensional crystals, most notably $\pi$-conjugated linear chains Dalton et al. (1987); Bredas et al. (1994), constitute prime examples to probe enhanced optical responses at non-linear orders. Because these crystals are centrosymmetric, i.e., they feature an inversion symmetry, their optical responses occur at odd orders in optical fields. The first- and third-order optical responses of one-dimensional $\pi$-conjugated chains have long been investigated, both experimentally  Fincher et al. (1979); Bredas et al. (1994) and computationally Hermann and Ducuing (1974); Rustagi and Ducuing (1974). Of central interest in this context is the third-order hyperpolarizability, i.e., the third-order electric susceptibility $\chi^{(3)}$ in the static limit of vanishing optical field frequency, $\omega\rightarrow 0$, which quantifies the non-linear polarization response, ${P}^{(3)}=\chi^{(3)}E_{x}^{3}$, to electric field $E_{x}$. In $\pi$-conjugated chains, $\chi^{(3)}$ is known to acquire colossal values. For example, $\chi^{(3)}$ was inferred to be as high as $10^{-18}$ m2/V2 in polyacetylene Boyd (2019), $10^{-19}$ m2/V2 in polydiacetylene, and $10^{-20}$ m2/V2 in polyarylenes Bredas et al. (1994). This unusually large hyperpolarizability has been heuristically ascribed to the responses of $p_{z}$-electrons forming the conjugated $\pi$-bonds. However, a rigorous and general theoretical understanding of the origin of these non-linear optical properties remains yet to be achieved.
@@
Figure 1: Crystalline topology and quantum geometrically-enhanced hyperpolarizability in one-dimensional crystals. (a) Schematic illustration of a one-dimensional inversion-symmetric model system with a topological invariant $\mathbb{Z}_{2}$ reflected by a Wannier charge center shift $\langle x\rangle$ and by the presence of a symmetry-protected exponentially-localized edge state, |$\ket{\psi_{\text{edge}}}$. The model system with unit cell parameter $a$ realizes a bipartite basis $A,B$, and effective hoppings $t_{1},t_{2}$ across bonds with lengths $a_{1},a_{2}$, and is coupled to an optical field $E_{x}(\omega)$ introduced by photons with energy $\hbar\omega$. The minimal spread of the Wannier function $\langle\Delta x^{2}\rangle$ is topologically-bounded, and evaluates to the band-averaged quantum metric $\langle g_{xx}\rangle$ Marzari and Vanderbilt (1997), which determines the non-linear polarizability. (b) Control of the band gap and band topology with lattice strain ($\lambda$) across a topological phase transition (TPT). (c) Discontinuous control of the non-linear optical properties, such as the time-dependent third-order polarization $P^{(3)}(t)$ from a three-photon coupling, across a strain-induced TPT.
@@
Figure 2: Quantum geometry and (non-)linear optics in one-dimensional crystals. (a) Quantum metric (red) over momentum-space as a function of the lattice strain parameter $(\lambda)$ realized in the effective low-energy Hamiltonian. The critical strain $(\lambda=\lambda_{c})$ induces a topological phase transition across which the metric scales up at the band edge, overcoming the minimal average metric value per BZ (blue) necessary for satisfying the topological bound. (b) Band-averaged metric $\langle g_{xx}\rangle$ as a function of strain and hopping amplitude ratio $t_{2}/t_{1}$. The corresponding optical absorptivity is manifestly larger and topologically lower-bounded in the topologically nontrivial phase. (c) Spectral resolution of strain-dependent optical conductivity $\sigma_{xx}(\omega)$ and cutoff ($\Lambda$) dependent optical weights $W_{xx}(\Lambda)=\int^{\Lambda}_{0}~{}\mathop{}\!\mathrm{d}\omega~{}\frac{\sigma_{xx}(\omega)}{\omega}$ (black) in a topological chain with hoppings ($t_{1}=2.15~{}\text{eV}$, $t_{2}=2.85~{}\text{eV}$) as a function of the bounding value for $\Lambda\rightarrow\infty$ (blue). (d) Lattice strain control of the topologically-bounded hyperpolarizability $\chi^{(3)}$. Under the relatively-high values of strain $\lambda/\lambda_{0}$, where $\lambda_{0}$ is a unit strain, the anomalous hyperpolarizability of the $\mathbb{Z}_{2}$-topological phase can be quenched. In the subcritical strain limit, we find that due to distinct realized quantum metrics, the $\chi^{(3)}$ of a trivial phase (red) is multiple times smaller than the $\chi^{(3)}$ of a topological phase (blue) with identical band energies. At vanishing strain, we note that the effective model fitted to the $t_{1}/t_{2}$ of polyacetylene yields $\chi^{(3)}=8$$\times$ 10-19 m2/V2, in excellent agreement with the corresponding experimental value, $\chi^{(3)}\approx$ 10-18 m2/V2 Fincher et al. (1979); Dalton et al. (1987); Bredas et al. (1994).
@@
In this Letter, we bridge this gap in knowledge by deploying the modern notions of band topology and quantum geometry. We show that the crystalline-symmetry-protected topology of $\pi$-conjugated chains induces a lower bound on the quantum metric, and thus on the hyperpolarizability. Then, we use numerical simulations to demonstrate the tunability of non-linear, quantum geometry-driven optical responses of a variety of one-dimensional crystals in which the band topology is controllable by an external parameter, such as lattice strain. Finally, we propose a general semiclassical picture to explain these remarkable optical properties, based on the scaling of quantum metric and band gaps with external parameters. Overall, our findings provide a novel interpretation of otherwise elusive experimental observations, along with guiding principles for identifying and engineering materials for non-linear optics.
@@
We now elucidate how, in one-dimensional (1D) crystals, the crystalline topology gives rise to quantum geometry and, as a direct consequence, enhances their optical responses. We define a non-Abelian multiband quantum geometric tensor, ${(\textbf{Q}^{m}_{ij})_{np}=\braket{\partial_{k_{i}}u_{n{\bf k}}}{u_{m{\bf k}}}\braket{u_{m{\bf k}}}{\partial_{k_{j}}u_{p{\bf k}}}}$ over momentum space, with its real part, commonly known as the non-Abelian quantum metric, being $(\textbf{g}^{m}_{ij})_{np}=\text{Re}~{}(\textbf{Q}^{m}_{ij})_{np}$ Provost and Vallee (1980); Ma et al. (2010); Bouhon et al. (2023). The multiband quantum metric encodes the optical transition dipole matrix elements, $(\textbf{g}^{m}_{xx})_{nn}=|\bra{\psi_{n{\bf k}}}\hat{x}\ket{\psi_{m{\bf k}}}|^{2}$ Ahn et al. (2020, 2021), which is at the heart of both linear and non-linear responses. The multiband quantum metric of electrons realizes lower bounds due to 1D crystalline topologies that translate into minimal values of (higher-order) optical susceptibilities,
$\langle g_{xx}\rangle\equiv\sum^{\text{unocc}}_{m}\sum_{k}\text{Tr}~{}{\textbf{g}}^{m}_{xx}\geq\frac{a^{2}}{4}\nu_{\text{FK}},$ (1)
where $\langle g_{xx}\rangle$ is band-averaged quantum metric and ${\nu_{\text{FK}}\in\mathbb{Z}_{2}}$ the Fu-Kane parity invariant protected by the $\mathcal{I}$-symmetry Fu and Kane (2007). In the context of the maximally localized Wannier functions, $\langle g_{xx}\rangle=\langle{\Delta x^{2}}\rangle$ Marzari and Vanderbilt (1997). Hence, the invariant imposes a lower-bound constraint on the spread of the wavefunction in the real-space basis. The derivation of Eq. (1) is provided in Sec. I of the Supplemental Material (SM). In topological crystalline phases, any topologically nontrivial band realizes $\nu_{\text{FK}}=1$. In the following, we demonstrate that this bound yields a topological enhancement of the multiband geometry-dependent response functions within the first and third-order optical responses.
@@
$\sigma_{xx}(\omega)=\frac{e^{2}}{2\hbar^{2}}\int_{\text{BZ}}\frac{\mathop{}\!\mathrm{d}k}{2\pi}\sum_{n,m}~{}f_{nm}E_{mn}\text{Tr}~{}{\textbf{g}}^{m}_{xx}\delta(\omega-E_{mn}/\hbar),$ (2)
where $f_{nm}=f_{n}-f_{m}$ are the thermal Fermi-Dirac occupation factors and $E_{mn}$ is the difference in energy between valence and conduction bands $n$ and $m$, respectively. For crystals possessing the same band structure, the distinct first order optical conductivity is purely determined by the multiband quantum metric. This allows one to uniquely distinguish band topologies by inducing the corresponding band geometries. Importantly, changes in band topology, as realized through external parameters such as lattice strain, as shown in Fig. 1(a,b), result in changes in the optical response. As we shall discuss in the representative case of $\pi$-conjugate chains and related one-dimensional crystals, lattice strain can be leveraged to engineer the opacity and optical non-linearities, as illustrated in Fig. 1(c).
@@
$\chi^{(3)}=\frac{e^{4}}{\epsilon_{0}}\int_{\text{BZ}}\frac{\mathop{}\!\mathrm{d}k}{2\pi}~{}\sum_{n,m,p,q}\frac{(\textbf{g}^{p}_{xx})_{nm}(\textbf{g}^{q}_{xx})_{mn}}{E_{mn}E_{pn}E_{qn}},$ (3)
@@
where $\alpha^{(3)}=\frac{1}{2}\pi^{3}e^{4}m_{e}^{3}\epsilon^{-1}_{0}\hbar^{-6}=6\times 10^{50}~{}\text{m}^{-5}/\text{V}^{2}$, $m_{e}$ is the electron mass, and $n$ is the number density of electrons in the crystal. The derivation of Eq. (4) is provided in Sec. IV of the SM SI .
From models to real materials.— We next develop an effective, two-band Hamiltonian to capture the geometry induced by $\mathbb{Z}_{2}$ obstructed topology in the presence of lattice strain. This model is schematically depicted in Fig. 1(a). The Hamiltonian that minimally realizes such general phenomenology on a quasi-one-dimensional chain takes the form of a strained SSH model Su et al. (1979, 1980),
@@
where $\lambda>0$ is an external tensile stress parameter, ${a(\lambda)\equiv a_{1}(\lambda)+a_{2}(\lambda)}$ is the strain-renormalized lattice parameter with bond lengths $a_{j}(\lambda)=a_{j}(0)[1+\lambda]$, and ${t_{j}(\lambda)=t_{j}(0)e^{-\gamma_{j}[a_{j}(\lambda)-a_{j}(0)]}}$ are the strain-dependent hopping amplitudes, where $\gamma_{j}$’s parameterize bond stiffness Tepliakov et al. (2023). We validate our strain-dependent models with hopping parameters obtained from first-principles calculations Jankowski et al. (2024). We correspondingly present our numerical results in Fig. 2. In Fig. 2(a), we show how the profile of quantum metric $g_{xx}$ over $k$-space, with dominant contributions at $k=\pi/a$ at the band inversion point, changes with the strain $\lambda$. Upon crossing the critical strain $\lambda=\lambda_{c}$, the band edge metric contributions provides for an average $\langle g_{xx}\rangle$ that satisfies the bound in the topological phase ($\nu_{\text{FK}}=1$). Fig. 2(b) presents a dependence of the optical weight, i.e., weighted integrated optical conductivity, or equivalently, integrated interband contribution to the imaginary part of dielectric constant ${W_{xx}=\int^{\infty}_{0}~{}\mathop{}\!\mathrm{d}\omega~{}\frac{\sigma_{xx}(\omega)}{\omega}=\frac{\pi e^{2}}{\hbar}\langle g_{xx}\rangle}$, which is also lower-bounded by the topological invariant. In. Fig. 2(c), we benchmark the first-order optical conductivity and show how its contributions to the bounded optical weights gradually arise. Fig. 2(d) shows how the colossal values of $\chi^{(3)}$ arise with strain $\lambda$, before vanishing sharply upon transitioning to the trivial phase across the critical point. We remark that the value of $\chi^{(3)}$ at vanishing $\lambda$ is in excellent agreement with experimental observations. Further details on the model-dependent parameterization and controllability of quantum geometry within $\pi$-conjugated chains are given in Sec. III of SM SI .
@@
Figure 3:  Examples of materials for crystalline topology-induced, quantum-geometrically enhanced non-linear optics. From top to bottom: polyacetylene, poly-($p$-phenylene), polypentacene, and graphene nanoribbons. Under external control, such as lattice strains, these materials can realize $\mathcal{I}$-symmetry protected nontrivial $\mathbb{Z}_{2}$ topology and topologically-enhanced hyperpolarizability $\chi^{(3)}$, which culminates in non-linear polarization $P^{(3)}$. Further chemical functionalizations of such $\mathbb{Z}_{2}$-topological materials, e.g., through perhalogenations, can provide a further knob to tune the quantum metric and higher-order susceptibilities.
@@
We stress that this model Hamiltonian and accompanying results are effectively realized by the low-energy bands of an ample variety of one-dimensional $\pi$-conjugate systems, such as polyacetylene Su et al. (1979, 1980), polyacenes Cirera et al. (2020), polyphenylenes, and graphene nanoribbons Tepliakov et al. (2023), as overviewed in Fig. 3. Their low-energy valence and conduction bands are universally governed by an effective physics of hopping between the Wannier orbitals $A,B$ that are predominantly constituted by the weighted combinations of the atomic orbitals with $p_{z}$ character Su et al. (1979); Cirera et al. (2020); Jankowski et al. (2024). Indeed, the introduction of additional bands does not alter out conclusions, as we demonstrate in Sec. V of the SM SI , where we also provide an extended discussion of multiband models. The reason for this traces back to the fact that the additional bands will experience higher energy gaps $E_{mn}$, suppressing the $\chi^{(3)}$ contributions as $1/E^{3}_{mn}$. Furthermore, the metric similarly scales with $1/E^{2}_{mn}$ [cf. Fig. 2(a)], which provides only a marginal contribution to the numerator of $\chi^{(3)}$.
@@
From quantum to semiclassical picture.— We now present an intuitive picture of the topologically-induced quantum geometric phenomenology. Semiclassically, electrons in band gap insulators can be modeled with an effective non-linear anharmonic oscillator Hamiltonian, ${H=\frac{p^{2}}{2m}+U(x)}$ with a $\mathcal{I}$-symmetric confining non-linear potential ${U(x)=\frac{1}{2}\kappa x^{2}+\frac{1}{4}bx^{4}}$, where $\kappa$ is an effective spring constant, $\kappa=m\omega_{0}^{2}$, in which $\omega_{0}$ is a resonant frequency and $b$ quantifies the degree of anharmonicity. Notably, ${\kappa~{}\propto 1/\langle g_{xx}\rangle=1/\langle\Delta x^{2}\rangle}$ Onishi and Fu (2024), where $\langle\Delta x^{2}\rangle$ is a spatial spread of a localized state wavefunction, as representable in terms of Wannier functions Marzari and Vanderbilt (1997). From the theory of non-linear anharmonic oscillator discussed in Sec. VI of the SM SI , one expects $\chi^{(3)}\propto b/\kappa^{4}$, by recognizing that, typically, $b\propto\kappa/\langle g_{xx}\rangle$, explaining the effective scaling of the ${\chi^{(3)}\propto\langle g_{xx}\rangle^{2}}$ within a full quantum-mechanical treatment. The semiclassical picture, in connection to quantum geometry, further supports the interpretation of our key result. Namely, the topological symmetry-indicated winding of the Bloch states in a one-dimensional insulator: (i) induces a geometric spread in the localization of the bounded electrons, which (ii) semiclassically reflects an effective spring constant and its inteplay with anharmonicity. The latter determine the non-linear optical response of the driven localized electrons within an anharmonic oscillator picture.

### v2

@@ Abstract
We demonstrate that higher-order electric susceptibilities in crystals can be enhanced and understood through nontrivial topological invariants and quantum geometry, using one-dimensional $\pi$-conjugated chains as representative model systems. First, we show that the crystalline-symmetry-protected topology of these chains imposes a lower bound on their quantum metric and hyperpolarizabilities. Second, we employ numerical simulations to reveal the tunability of nonlinear, quantum geometry-driven optical responses in various one-dimensional crystals in which band topology can be externally controlled. Third, we develop a semiclassical picture to deliver an intuitive understanding of these effects. Our findings offer a firm interpretation of otherwise elusive experimental observations of colossal hyperpolarizabilities and establish guidelines for designing topological materials of any dimensionality with enhanced nonlinear optical properties.
@@
Introduction.— Nonlinear optics offers a rich platform to explore a broad range of intriguing phenomena, with implications ranging from fundamental science to engineering Boyd (2019). One-dimensional crystals, most notably $\pi$-conjugated linear chains Agrawal et al. (1978); Dalton et al. (1987); Bredas et al. (1994), constitute prime examples to probe enhanced optical responses at nonlinear orders. Because these crystals are centrosymmetric, i.e., they feature an inversion symmetry, their optical responses occur at odd orders in optical fields. The first- and third-order optical responses of one-dimensional $\pi$-conjugated chains have long been investigated, both experimentally  Fincher et al. (1979); Beratan et al. (1987); Sinclair et al. (1988); Samuel et al. (1994); Ledoux et al. (1999); Bredas et al. (1994) and computationally Hermann and Ducuing (1974); Rustagi and Ducuing (1974); Soos and Hayden (1989); Villesuzanne et al. (1992); Huang et al. (1992); Shuai and Brédas (1992, 1992); Schmidt and Springborg (1999); Hu et al. (2007); Lacivita et al. (2012); Luzanov (2023). Of central interest in this context is the third-order hyperpolarizability, i.e., the third-order electric susceptibility $\chi^{(3)}$ in the static limit of vanishing optical field frequency, $\omega\rightarrow 0$, which quantifies the nonlinear polarization response, ${P}^{(3)}=\chi^{(3)}E_{x}^{3}$, to electric field $E_{x}$ Pronin et al. (1995). In $\pi$-conjugated chains, $\chi^{(3)}$ is known to acquire colossal values. For example, $\chi^{(3)}$ was inferred to be as high as $10^{-18}$ m2/V2 in polyacetylene Boyd (2019), $10^{-19}$ m2/V2 in polydiacetylene, and $10^{-20}$ m2/V2 in polyarylenes Bredas et al. (1994). This unusually large hyperpolarizability has been heuristically ascribed to the responses of $p_{z}$-electrons forming the conjugated $\pi$-bonds Boyd (2019). However, a general theoretical understanding of the phenomenological origin of these nonlinear optical properties remains yet to be achieved Yu et al. (1989a); Hagler and Heeger (1994); Takahashi (1995); Yaron (1996); Takahashi (1997a, b); Matsuzaki et al. (1998); Margulis and Gaiduk (1998); Jiang and Xu (2005); Xu and Jiang (2005); Díaz-Ponce (2017).
@@
Figure 1: Crystalline topology and quantum geometrically-enhanced hyperpolarizability in one-dimensional crystals. (a) Schematic illustration of a one-dimensional inversion-symmetric model system with a topological invariant $\mathbb{Z}_{2}$ reflected by a Wannier charge center shift $\langle x\rangle$ and by the presence of a symmetry-protected exponentially-localized edge state, |$\ket{\psi_{\text{edge}}}$. The model system with unit cell parameter $a$ realizes a bipartite basis $A,B$, and effective hoppings $t_{1},t_{2}$ across bonds with lengths $a_{1},a_{2}$, and is coupled to an optical field $E_{x}(\omega)$ introduced by photons with energy $\hbar\omega$. The minimal spread of the Wannier function $\langle\Delta x^{2}\rangle$ is topologically-bounded, and evaluates to the band-averaged quantum metric $\langle g_{xx}\rangle$ Marzari and Vanderbilt (1997), which determines the nonlinear polarizability. (b) Control of the band gap and band topology with lattice strain ($\lambda$) across a topological phase transition (TPT). (c) Discontinuous control of the nonlinear optical properties, such as the time-dependent third-order polarization $P^{(3)}(t)$ from a three-photon coupling, across a strain-induced TPT.
@@
Figure 2: Quantum geometry and (non-)linear optics in one-dimensional crystals. (a) Quantum metric (red) over momentum-space as a function of the lattice strain parameter $(\lambda)$ realized in the effective low-energy Hamiltonian, Eq. (5), with hoppings $t_{1}(0)=2.15~{}\text{eV}$, $t_{2}(0)=2.85~{}\text{eV}$ [$t_{2}(0)/t_{1}(0)=1.33$]. The critical strain $(\lambda=\lambda_{c})$ induces a topological phase transition across which the metric scales up at the band edge at $k=\pm\pi/a(\lambda)$, overcoming the minimal average metric value per BZ (blue) necessary for satisfying the topological bound. (b) Band-averaged metric $\langle g_{xx}\rangle$ as a function of strain and hopping amplitude ratio $t_{2}(0)/t_{1}(0)$. The horizontal line corresponds to the chain with $t_{2}(0)/t_{1}(0)=1.33$. The corresponding optical absorptivity is manifestly larger and topologically lower-bounded in the topologically nontrivial phase. (c) Spectral resolution of strain-dependent optical conductivity $\sigma_{xx}(\omega)$ and cutoff ($\Lambda$) dependent optical weights $W_{xx}(\Lambda)=\int^{\Lambda}_{0}~{}\mathop{}\!\mathrm{d}\omega~{}\frac{\sigma_{xx}(\omega)}{\omega}$ (black), as a function of the bounding value for $\Lambda\rightarrow\infty$ (blue) in a topological chain with $t_{2}/t_{1}=1.33$. (d) Lattice strain control of the topologically-bounded hyperpolarizability $\chi^{(3)}$. Under the relatively-high values of strain $\lambda/\lambda_{0}$, where $\lambda_{0}$ is a unit strain, the anomalous hyperpolarizability of the $\mathbb{Z}_{2}$-topological phase can be quenched. In the subcritical strain limit, we find that due to distinct realized quantum metrics, the $\chi^{(3)}$ of a trivial phase (red) is multiple times smaller than the $\chi^{(3)}$ of a topological phase (blue) with identical band energies. At vanishing strain, we note that the effective model fitted to the $t_{2}/t_{1}$ of polyacetylene yields $\chi^{(3)}=8$$\times$ 10-19 m2/V2, in excellent agreement with the corresponding experimental value, $\chi^{(3)}\approx$ 10-18 m2/V2 Fincher et al. (1979); Dalton et al. (1987); Bredas et al. (1994).
@@
In this Letter, we bridge this gap by deploying the modern notions of band topology and quantum geometry. We show that the crystalline-symmetry-protected topology of $\pi$-conjugated chains induces a lower bound on the quantum metric, and thus on the hyperpolarizability. Then, we use numerical simulations to demonstrate the tunability of nonlinear, quantum geometry-driven optical responses of a variety of one-dimensional crystals in which the band topology is controllable by an external parameter, such as lattice strain. Finally, we propose a general semiclassical picture to explain these remarkable optical properties, based on the scaling of quantum metric and band gaps with external parameters. Importantly, our theory offers a topological explanation for the distinct hyperpolarizabilities of the cis and trans isomers of $\pi$-conjugated linear polymers Sinclair et al. (1988). Unlike the cis configuration, the topological windings of Bloch eigenvector in the trans configuration support the geometric enhancements of the optical matrix elements associated with the nonlinear optical responses. Overall, our findings provide a novel interpretation of otherwise elusive experimental observations, along with guiding principles for identifying and engineering materials for nonlinear optics.
@@
We now elucidate how, in one-dimensional (1D) crystals, the crystalline topology gives rise to quantum geometry and, as a direct consequence, enhances their optical responses. We define a non-Abelian multiband quantum geometric tensor, ${[\textbf{Q}^{(m)}_{ij}]_{np}=\braket{\partial_{k_{i}}u_{n{\bf k}}}{u_{m{\bf k}}}\braket{u_{m{\bf k}}}{\partial_{k_{j}}u_{p{\bf k}}}}$ over momentum space, with its real part, commonly known as the non-Abelian quantum metric, being $[\textbf{g}^{(m)}_{ij}]_{np}=\text{Re}~{}[\textbf{Q}^{(m)}_{ij}]_{np}$ Provost and Vallee (1980); Ma et al. (2010); Bouhon et al. (2023). The multiband quantum metric encodes the optical transition dipole matrix elements, $[\textbf{g}^{(m)}_{xx}]_{nn}=|\bra{\psi_{n{\bf k}}}\hat{x}\ket{\psi_{m{\bf k}}}|^{2}$ Ahn et al. (2020, 2021), which is at the heart of both linear and nonlinear responses. The multiband quantum metric of electrons realizes lower bounds due to 1D crystalline topologies that translate into minimal values of (higher-order) optical susceptibilities,
$\langle g_{xx}\rangle\equiv\sum^{\text{unocc}}_{m}\sum_{k}\text{Tr}~{}{\textbf{g}}^{(m)}_{xx}\geq\frac{a^{2}}{4}\nu_{\text{FK}},$ (1)
where $\langle g_{xx}\rangle$ is band-averaged quantum metric and ${\nu_{\text{FK}}\in\mathbb{Z}_{2}}$ the Fu-Kane parity invariant protected by the $\mathcal{I}$-symmetry Fu and Kane (2007). The bound can be derived from (i) non-trivial inversion symmetry indicators pinpointing how the irreducible representations of the obstructed bands transform Fu and Kane (2007); Kruthoff et al. (2017); Po et al. (2017); Bradlyn et al. (2017), (ii) translating these into real-space projectors Herzog-Arbeitman et al. (2022), and (iii) using an optimization argument of Ref. Herzog-Arbeitman et al. (2022). The technical details of the derivation are given in Sec. I of the Supplemental Material (SM) SI . In the context of the maximally localized Wannier functions, $\langle g_{xx}\rangle=\langle{\Delta x^{2}}\rangle$ Marzari and Vanderbilt (1997). Hence, the invariant imposes a lower-bound constraint on the spread of the wavefunction in the real-space basis. In topological crystalline phases, any topologically nontrivial band realizes $\nu_{\text{FK}}=1$. In the following, we demonstrate that this bound yields a topological enhancement of the multiband geometry-dependent response functions within the first and third-order optical responses.
@@
$\sigma_{xx}(\omega)=\frac{e^{2}}{2\hbar^{2}}\int_{\text{BZ}}\frac{\mathop{}\!\mathrm{d}k}{2\pi}\sum_{n,m}~{}f_{nm}E_{mn}\text{Tr}~{}{\textbf{g}}^{(m)}_{xx}\delta(\omega-E_{mn}/\hbar),$ (2)
where $f_{nm}=f_{n}-f_{m}$ are the thermal Fermi-Dirac occupation factors and $E_{mn}$ is the difference in energy between valence and conduction bands $n$ and $m$, respectively. For crystals possessing the same band structure, the distinct first order optical conductivity is purely determined by the multiband quantum metric. This allows one to uniquely distinguish band topologies by inducing the corresponding band geometries. Importantly, changes in band topology, as realized through external parameters such as lattice strain, as shown in Fig. 1(a,b), result in changes in the optical response. As we shall discuss in the representative case of $\pi$-conjugate chains and related one-dimensional crystals, lattice strain can be leveraged to engineer the opacity and optical nonlinearities, as illustrated in Fig. 1(c).
@@
$\chi^{(3)}=\frac{e^{4}}{\epsilon_{0}}\int_{\text{BZ}}\frac{\mathop{}\!\mathrm{d}k}{2\pi}~{}\sum_{n,m,p,q}\frac{[\textbf{g}^{(p)}_{xx}]_{nm}[\textbf{g}^{(q)}_{xx}]_{mn}}{E_{mn}E_{pn}E_{qn}},$ (3)
@@
where $\alpha^{(3)}=\frac{1}{2}\pi^{3}e^{4}m_{e}^{3}\epsilon^{-1}_{0}\hbar^{-6}=6\times 10^{50}~{}\text{m}^{-5}/\text{V}^{2}$, $m_{e}$ is the electron mass, and $n$ is the number density of electrons in the crystal. The derivation of Eq. (4) employs (i) the energy gap bound of Ref. Kivelson (1982) due to quantum metric (ii) combined with the sum rules of Ref. Souza et al. (2000), as employed in the linear susceptibility bound of Ref. Onishi and Fu (2024). The technical details of the derivation are given in Sec. IV of the SM SI .
From models to real materials.— We next employ an effective, two-band Hamiltonian to capture the geometry induced by $\mathbb{Z}_{2}$ obstructed topology in the presence of lattice strain. This model is schematically depicted in Fig. 1(a). The Hamiltonian that minimally realizes such general phenomenology on a quasi-one-dimensional chain takes the form of a strained SSH model Su et al. (1979, 1980),
@@
where $\lambda>0$ is an external tensile stress parameter, ${a(\lambda)\equiv a_{1}(\lambda)+a_{2}(\lambda)}$ is the strain-renormalized lattice parameter with bond lengths $a_{j}(\lambda)=a_{j}(0)[1+\lambda]$, and ${t_{j}(\lambda)=t_{j}(0)e^{-\gamma_{j}[a_{j}(\lambda)-a_{j}(0)]}}$ are the strain-dependent hopping amplitudes, where $\gamma_{j}$’s parameterize bond stiffness Tepliakov et al. (2023). We validate our strain-dependent models with hopping parameters obtained from first-principles calculations Jankowski et al. (2025). We correspondingly present our numerical results in Fig. 2. In Fig. 2(a), we show how the profile of quantum metric $g_{xx}$ over $k$-space changes with the strain $\lambda$. The metric supporting the topological lower bound is enhanced at the band edge $k=\pi/a$, as the dominant contributions arise at the band inversion region, where the local derivatives of the Bloch eigenvectors winding across the BZ are maximized. Upon crossing the critical strain $\lambda=\lambda_{c}$, the band edge metric contributions provides for an average $\langle g_{xx}\rangle$ that satisfies the bound in the topological phase ($\nu_{\text{FK}}=1$). Fig. 2(b) presents a dependence of the optical weight, i.e., weighted integrated optical conductivity, or equivalently, integrated interband contribution to the imaginary part of dielectric constant ${W_{xx}=\int^{\infty}_{0}~{}\mathop{}\!\mathrm{d}\omega~{}\frac{\sigma_{xx}(\omega)}{\omega}=\frac{\pi e^{2}}{\hbar}\langle g_{xx}\rangle}$, which is also lower-bounded by the topological invariant. In. Fig. 2(c), we benchmark the first-order optical conductivity and show how its contributions to the bounded optical weights gradually arise. Fig. 2(d) shows how the colossal values of $\chi^{(3)}$ arise with strain $\lambda$, before vanishing sharply upon transitioning to the trivial phase across the critical point. We remark that the value of $\chi^{(3)}$ at vanishing $\lambda$ is in excellent agreement with experimental observations. Importantly, the bound of the quantum metric, as manifested in Fig. 2(c-d), provides for the robustness of lower-bounded optical responses of topological phases, especially away from the singular critical points that correspond to the strain-induced phase transitions. Further details on the model-dependent parameterization and controllability of quantum geometry within $\pi$-conjugated chains are given in Sec. III of SM SI .
@@
Figure 3:  Examples of materials for crystalline topology-induced, quantum-geometrically enhanced nonlinear optics. From top to bottom: polyacetylene, poly-($p$-phenylene), polypentacene, and graphene nanoribbons. Under external control, such as lattice strains, these materials can realize $\mathcal{I}$-symmetry protected nontrivial $\mathbb{Z}_{2}$ topology and topologically-enhanced hyperpolarizability $\chi^{(3)}$, which culminates in nonlinear polarization $P^{(3)}$. Further chemical functionalizations of such $\mathbb{Z}_{2}$-topological materials, e.g., through perhalogenations, can provide a further knob to tune the quantum metric and higher-order susceptibilities.
@@
We stress that this model Hamiltonian and accompanying results are effectively realized by the low-energy bands of an ample variety of one-dimensional $\pi$-conjugate systems, such as polyacetylene Su et al. (1979, 1980), polyacenes Cirera et al. (2020), polyphenylenes, and graphene nanoribbons Tepliakov et al. (2023), as overviewed in Fig. 3. Their low-energy valence and conduction bands are universally governed by an effective physics of hopping between the Wannier orbitals $A,B$ that are predominantly constituted by the weighted combinations of the atomic orbitals with $p_{z}$ character Su et al. (1979); Cirera et al. (2020); Jankowski et al. (2025). The physical implication of our findings is that the materials shown in Fig. 3 are expected to exhibit quantum-geometrically enhanced hyperpolarizabilities that originate from their nontrivial, inversion symmetry-protected band topology. Indeed, the introduction of additional bands does not alter out conclusions, as we demonstrate in Sec. V of the SM SI , where we also provide an extended discussion of multiband models. The reason for this traces back to the fact that the additional bands will experience higher energy gaps $E_{mn}$, suppressing the $\chi^{(3)}$ contributions as $1/E^{3}_{mn}$. Furthermore, the metric similarly scales with $1/E^{2}_{mn}$ [cf. Fig. 2(a)], which for topologically trivial additional bands provides only a marginal contribution to the numerator of $\chi^{(3)}$.
@@
From quantum to semiclassical picture.— We now present an intuitive picture of the topologically-induced quantum geometric phenomenology. Semiclassically, electrons in band gap insulators can be modeled with an effective nonlinear anharmonic oscillator Hamiltonian, ${H=\frac{p^{2}}{2m_{e}}+U(x)}$ with a $\mathcal{I}$-symmetric confining nonlinear potential ${U(x)=\frac{1}{2}\kappa x^{2}+\frac{1}{4}bx^{4}}$, where $\kappa$ is an effective spring constant, $\kappa=m_{e}\omega_{0}^{2}$, in which $\omega_{0}$ is a resonant frequency and $b$ quantifies the degree of anharmonicity. Notably, ${\kappa~{}\propto 1/\langle g_{xx}\rangle=1/\langle\Delta x^{2}\rangle}$ Onishi and Fu (2024), where $\langle\Delta x^{2}\rangle$ is a spatial spread of a localized state wavefunction, as representable in terms of Wannier functions Marzari and Vanderbilt (1997). From the theory of nonlinear anharmonic oscillator discussed in Sec. VI of the SM SI , one expects $\chi^{(3)}\propto b/\kappa^{4}$, by recognizing that, typically, $b\propto\kappa/\langle g_{xx}\rangle$, explaining the effective scaling of the ${\chi^{(3)}\propto\langle g_{xx}\rangle^{2}}$ within a full quantum-mechanical treatment. The semiclassical picture, in connection to quantum geometry, further supports the interpretation of our key result. Namely, the topological symmetry-indicated winding of the Bloch states in a one-dimensional insulator: (i) induces a geometric spread in the localization of the bounded electrons, which (ii) semiclassically reflects an effective spring constant and its inteplay with anharmonicity. The latter determine the nonlinear optical response of the driven localized electrons within an anharmonic oscillator picture.